OpenStudy (bananas1234):

Arrange these numbers from least to greatest: 3/5 , 5/7 , 2/9

OpenStudy (bananas1234):

my options are: 2/9, 3/5, 5/7 2/9, 5/7, 3/5 5/7, 3/5, 2/9 3/5, 2/9, 5/7

OpenStudy (triciaal):

one approach more equivalent fractions find the LCD of 5, 7 and 9 find the appropriate equivalents then place on the number line list from least to greatest

OpenStudy (anonymous):

It is the third one or C

OpenStudy (bananas1234):

Would it be the second one? @triciaal

OpenStudy (mathstudent55):

Use a calculator to find the decimal equivalents. Then you can compare quickly.

OpenStudy (bananas1234):

Thats a good way to do it. @mathstudent55

OpenStudy (triciaal):

did not do it

OpenStudy (triciaal):

what is the lcd?

OpenStudy (bananas1234):

wait....would it be A?

OpenStudy (bananas1234):

3/5 = 189/315 5/7 = 225/315 2/9 = 70/315 so the least is 2/9 then 3/5 and the 5/7

OpenStudy (mathstudent55):

3 is a little more than half of 5, so 3/5 is more than 1/2. 5 is more than half of 7, so 5/7 is more than 1/2. 2 is much less than half of 9, so 2/9 is much less than 1/2. For sure 2/9 is the smallest fraction. Now just compare 3/5 to 5/7.

OpenStudy (mathstudent55):